# Zhaoxiang Zhang9, Yuanxing Zhang10, Ge Zhang3,7, Ke Xu1, Jie Fu3,11*, Wenhao Huang2,3

### v2

@@ 2.3 Supervised Fine-Tuning
As shown in Table 7, our model undergoes comprehensive and systematic supervised fine-tuning (SFT) with 16 different tasks and 34 diverse open-source datasets. The chat template used for SFT is the same as that used for Yi-6B-Chat [47], and only the assistant responses are supervised. We refer the reader to Appendix D for more details about the hyperparameters and prompt templates.
@@ 3 Experiments
In this section, we present our quantitative evaluation results across various domains: image-related tasks (§3.1), speech-related tasks (§3.2), and video-related tasks (§3.3). Due to the lack of benchmarks for several advanced and emergent abilities of any-to-any multimodal LLMs, we also provide numerous qualitative demonstrations (§3.4) to demonstrate these capabilities. We refer the reader to Appendix E for more details, including the decoding hyperparameters and prompt templates.
@@ 3.1 Image-Related Tasks
We compare our models with Emu [18], SEED-LLaMA [20], AnyGPT [21], Flamingo [7], Kosmos-1 [51], MetaLM [52], IDEFICS [53], CM3Leon [22], and InstructBLIP [54]. We evaluate our models in diverse tasks, including: (1) image captioning on MS-COCO [55] Karpathy test split with CIDEr score [56] as the metric, (2) three visual question-answering benchmarks, i.e., VQAv2 [57] (test-dev split), OK-VQA [58] (val split), and VizWiz [59], with VQA accuracy as the metric, and (3) SEED-Bench [60], a comprehensive visual question-answering benchmark including 9 dimensions with MCQ accuracy as the metric. The scores for all baselines are copied from their reports. As shown in Table 3, our MIO-Instruct is ranked in the top group among all baselines, demonstrating its competitive image understanding performance. Although SEED-LLaMA achieved better scores compared to our model, we additionally support the speech modality. It is also noteworthy that MIO, with a size of approximately 7 billion parameters, outperforms several larger models such as Emu-14B and even IDEFICS-80B.
@@
We compare our models with Emu [18], SEED-LLaMA [20], GILL [61], and AnyGPT [21] for image generation. We use two benchmarks, i.e., MS-COCO [55] Karpathy test split and Flickr30K [62]. Following GILL [61] and SEED-LLaMA [20], we use CLIP-I as the metric that evaluates the similarity between the generated images and the ground-truth images with the image encoder in CLIP [63]. As shown in Table 4, the pre-trained model and instruction-tuned model of MIO both have competitive image generation capabilities. Note that beyond single image generation abilities, our model can also exhibit multi-image generation capabilities such as generating visual stories, image sequences, and even visual thoughts as illustrated in §3.4.
@@ 3.2 Speech-Related Tasks
We evaluate the speech understanding and generation abilities of MIO on ASR and TTS tasks. Wav2vec 2.0 [64], Whisper Large V2 [65], and AnyGPT [21] are the baselines for ASR tasks, while VALL-E [66], USLM [42] , and AnyGPT [21] are the baselines for TTS tasks. The test set used for ASR evaluation is LibriSpeech [67], while the test set used for TTS evaluation is VCTK [68] following AnyGPT [21]’s practice. The Whisper medium model is used to transcribe the speech generated for the TTS task. The WER (word error rate) is computed by comparing the generated transcribed text with the ground-truth transcription after text normalization [^4] https://github.com/openai/whisper/blob/main/whisper/normalizers/english.py. As shown in Table 3.2, our models exhibit speech performance comparable to the speech-specific baselines and outperform the AnyGPT baseline. It is important to note that although AnyGPT is capable of generating content tokens for speech, it lacks the ability to generate timbre tokens, which necessitates the use of an additional voice cloning model. In contrast, our models generate both content and timbre tokens, making the TTS tasks more challenging for our models compared to AnyGPT. Nonetheless, after instruction tuning, our model still achieves better TTS performance.
@@ 3.3 Video-Related Tasks
We compare MIO with Flamingo [7], BLIP-2 [6], InstructBLIP [54], Emu [18], and SEED-LLaMA [20] for video understanding. The models are evaluated on the MSVDQA [69] and MSRVTT-QA [70]. The results are presented in Table 6. Our model achieves the highest scores compared to all baselines. Due to the lack of video (frame sequence) generation benchmarks in our setting, we provide video generation examples in §3.4. These results demonstrate the superior performance of our models in both video understanding and video generation.
@@ 3.4 Demonstrations
Figure 2: Demonstrations for the basic abilities of MIO.
Figure 3: Demonstrations for the advanced abilities of MIO.
We illustrate the basic and advanced abilities of MIO in Figure 2 and 3. The basic abilities of MIO involve image understanding and generation, video understanding and generation, ASR, and TTS. The advanced abilities of MIO are based on its any-to-any and multimodal interleaved sequence generation features. These abilities involve visual storytelling (i.e., interleaved video-text generation), chain of visual thought, speech-in-speech-out, instructional image editing, visual guideline generation, etc. We refer the readers to Appendix F for more demonstrations including multimodal chain of thought and multimodal in-context learning.
@@ 4.1 Multimodal LLMs
With the rapid success of Large Language Models (LLMs), current multimodal LLMs (MM-LLMs) are typically built on a pre-trained LLM backbone and are endowed with the ability to understand multiple modalities [71, 72, 73, 74, 75, 76, 77]. Generally, these MM-LLMs align the representations of images obtained from visual encoders with the text embedding space, thereby leveraging the powerful capabilities of the foundational models. For example, BLIP-2 [6] uses CLIP-ViT [63] to extract high-level features from images and then employs a Q-Former to compress the number of image tokens and further align image tokens with the LLM embeddings. In contrast, LLaVA [17, 78] utilizes a simple linear projection or MLP as the connector between the image encoder and the LLM backbone. These models demonstrate strong multimodal understanding abilities, achieving significant progress in tasks such as visual question answering, visual commonsense reasoning, chart understanding, etc.
Additionally, beyond images, other MM-LLMs have also focused on modalities such as speech and video. For instance, LLaSM [79] and InternVideo [80, 81] are MM-LLMs designed for speech and video understanding, respectively. These models adopt a similar architectural design to BLIP-2 or LLaVA but redesign modality-specific encoders tailored for speech and video.
Recently, increasing attention has been given to unifying multiple modalities within a single MM-LLM. For example, ImageBind [82] develops encoders suited for multiple modalities such as images, videos, audio, heat maps, among others, while OmniBind [83] trains an omni-representation model by aligning encoders across four modalities: audio, language, images, and 3D objects.
@@ 4.2 Any-to-Any MM-LLMs
To enable multimodal generation in MM-LLMs, a straightforward approach is to allow these models to call external multimodal generation tools, such as Stable Diffusion [49] or text-to-speech (TTS) tools [37, 84, 1]. However, as highlighted in the Gemini technical report [30], relying on an intermediate natural language interface can limit the model’s ability to express images. If a model cannot natively output images, it will not be able to generate images with prompts of interleaved sequences of image and text. This claim is in line with our distinction between descriptive image generation and context-aware image generation, as discussed in §2.2.
As a result, recent works focus on the unification of multimodal understanding and generation in a single model (i.e., any-to-any MM-LLMs), enabling the generation of multimodal tokens without natural language as an interface. These models typically follow different approaches, depending on how images are represented in both input and output sides. For example, the Discrete-In-Discrete-Out (DIDO) approach has been explored in works such as SEED-LLaMA [20], AnyGPT [21], and Chameleon [23]. Continuous-In-Discrete-Out (CIDO) methods have been implemented in models like DaVinCi [29], Gemini [30], and Unified-IO 2 [36]. The Continuous-In-Continuous-Out (CICO) approach is used in models such as Emu [18, 19], DreamLLM [31], and MiniGPT-5 [32]. Another approach, the integration of autoregression and diffusion (AR + Diff), can be seen in models like Transfusion [25], Show-o [34], and [35]’s.
However, these models face specific limitations. DreamLLM (CICI,  [31]) and CIDO models suffer from inconsistencies between input and output forms for multimodal data, making it difficult for them to natively support the generation of interleaved multimodal sequences where an image functions in a coupled way as both input and output. Emu2 (CICO,  [19]) struggles with the challenges of the mean square error (MSE) loss used for training continuous output representations, as well as with the uni-modal assumption of the Gaussian distribution in the MSE loss. Transfusion (AR + Diff,  [25]) applies noise to images from the input side to support multimodal generation with diffusion modeling, and relies on VAE [85] features rather than CLIP [63] features for denoising, which largely trade off the multimodal understanding abilities.
To mitigate these issues, we adopt the DIDO approach in our work, despite the information loss caused by vector quantization. A comprehensive comparison of our models with other any-to-any MM-LLMs is presented in Table 1.
@@ 5 Conclusion
Appendix A Limitations
Our work has several limitations. First, due to the information loss caused by vector quantization procedures, our model is not applicable for OCR-related images, or other images with highly detailed features. This is also a vital issue faced by CM3Leon [22] and Chameleon [86]. Second, due to the lack of voice-controlled speech data, our model cannot control the generated timbre yet. Third, although our model can generate frame sequences as video representation, it cannot generate the raw video data which is in the continuous form. We leave these issues as the research questions for the future work.
Appendix B Pre-training Data
@@
    Image-text paired data: SBU [87], CC3M [88], LAION-COCO [89], and JourneyDB [90], where JourneyDB only serves for image generation.
@@
    Language-only data: RefinedWeb [91].
@@
    Image-text interleaved data: OBELICS [53], MMC4-core-ff [92].
@@
    Video-text paired data: WebVid-10M [93].
@@
    Video-text interleaved data: HowTo-100M [94], Youtube-Temporal-180M [95].
@@
    Speech-text paired data: Libriheavy [96].
@@
We have different data processing procedures for different data types illustrated in §B following Emu [18] and Qwen-VL [97]:
@@
    Language-only data: we use the same data processing pipeline as used in Yi [47].
@@
    Image-text interleaved data: we filter the data using a CLIP score threshold of 0.25, and follow the same procedure as illustrated in Emu [18].
@@
    Video-text interleaved data: We first use PySceneDetect to extract key frames from the video based on scene changes, following the practice of Stable Video Diffusion [98]. Then, for each video clip between two key frames, we extract a central frame for textual caption generation with BLIP-2 [6]. Additionally, the video clips between key frames are processed using ASR (automatic speech recognition) tools to extract subtitles. The ASR text and captions are then integrated and refined using Yi-34B-Chat [47], resulting in a single text segment. These text segments, along with the key frames and central frames, form the video-text interleaved data.
@@
Appendix C Pre-training Details

### v3

@@ 2.3 Supervised Fine-Tuning
As shown in Table 9, our model undergoes comprehensive and systematic supervised fine-tuning (SFT) with 16 different tasks and 34 diverse open-source datasets. The chat template used for SFT is the same as that used for Yi-6B-Chat (AI et al., 2024), and only the assistant responses are supervised. We refer the reader to Appendix C for more details about the hyperparameters and prompt templates.
@@ 3 Experiments
In this section, we present our quantitative evaluation results across various domains: image-related tasks (§3.1), speech-related tasks (§3.2), and video-related tasks (§3.3). Due to the lack of benchmarks for several advanced and emergent abilities of any-to-any multimodal LLMs, we also provide numerous qualitative demonstrations (§3.5) to demonstrate these capabilities. We refer the reader to Appendix D for more details, including the decoding hyperparameters and prompt templates.
@@ 3.1 Image-Related Tasks
We compare our models with Emu (Sun et al., 2023c), SEED-LLaMA (Ge et al., 2023b), AnyGPT (Zhan et al., 2024), Flamingo (Alayrac et al., 2022), Kosmos-1 (Huang et al., 2023), MetaLM (Hao et al., 2022), IDEFICS (Laurençon et al., 2023), CM3Leon (Yu et al., 2023), InstructBLIP (Dai et al., 2023), Qwen-VL-Chat (Bai et al., 2023), and LLaVA 1.5 (Liu et al., 2023a). We evaluate our models in diverse tasks, including: (1) image captioning on MS-COCO (Lin et al., 2014) Karpathy test split with CIDEr score (Vedantam et al., 2014) as the metric, (2) three visual question-answering benchmarks, i.e., VQAv2 (Goyal et al., 2016) (test-dev split), OK-VQA (Marino et al., 2019) (val split), and VizWiz (Gurari et al., 2018), with VQA accuracy as the metric, and (3) SEED-Bench (Li et al., 2023a), a comprehensive visual question-answering benchmark including 9 dimensions with MCQ accuracy as the metric. The scores for all baselines are copied from their reports. As shown in Table 3, our MIO-Instruct is ranked in the top group among all baselines, demonstrating its competitive image understanding performance. Although SEED-LLaMA achieved better scores compared to our model, we additionally support the speech modality. It is also noteworthy that MIO, with a size of approximately 7 billion parameters, outperforms several larger models such as Emu-14B and even IDEFICS-80B.
@@
We compare our models with Emu (Sun et al., 2023c), SEED-LLaMA (Ge et al., 2023b), GILL (Koh et al., 2023), and AnyGPT (Zhan et al., 2024) for image generation. We use two benchmarks, i.e., MS-COCO (Lin et al., 2014) Karpathy test split and Flickr30K (Plummer et al., 2015). Following GILL (Koh et al., 2023) and SEED-LLaMA (Ge et al., 2023b), we use CLIP-I as the metric that evaluates the similarity between the generated images and the ground-truth images with the image encoder in CLIP (Radford et al., 2021). As shown in Table 4 and Table 12 the pre-trained model and instruction-tuned model of MIO both have competitive image generation capabilities. Note that beyond single image generation abilities, our model can also exhibit multi-image generation capabilities such as generating visual stories, image sequences, and even visual thoughts as illustrated in §3.5.
@@ 3.2 Speech-Related Tasks
We evaluate the speech understanding and generation abilities of MIO on ASR and TTS tasks. Wav2vec 2.0 (Baevski et al., 2020), Whisper Large V2 (Radford et al., 2023), and AnyGPT (Zhan et al., 2024) are the baselines for ASR tasks, while VALL-E (Wang et al., 2023a), USLM (Zhang et al., 2023b) , and AnyGPT (Zhan et al., 2024) are the baselines for TTS tasks. The test set used for ASR evaluation is LibriSpeech (Panayotov et al., 2015), while the test set used for TTS evaluation is VCTK (Veaux et al., 2017) following AnyGPT (Zhan et al., 2024)’s practice. The Whisper medium model is used to transcribe the speech generated for the TTS task. The WER (word error rate) is computed by comparing the generated transcribed text with the ground-truth transcription after text normalization [^4] https://github.com/openai/whisper/blob/main/whisper/normalizers/english.py.
@@
As shown in Table 3.2, our models exhibit speech performance comparable to the speech-specific baselines and outperform the AnyGPT baseline. It is important to note that although AnyGPT is capable of generating content tokens for speech, it lacks the ability to generate timbre tokens, which necessitates the use of an additional voice cloning model. In contrast, our models generate both content and timbre tokens, making the TTS tasks more challenging for our models compared to AnyGPT. Nonetheless, after instruction tuning, our model still achieves better TTS performance. More evaluations of the TTS and Speech-to-Speech generation performance are provided in Appendix E.3 and E.2.
@@ 3.3 Video-Related Tasks
We compare MIO with Flamingo (Alayrac et al., 2022), BLIP-2 (Li et al., 2023b), InstructBLIP (Dai et al., 2023), Emu (Sun et al., 2023c), and SEED-LLaMA (Ge et al., 2023b) for video understanding. The models are evaluated on the MSVDQA (Chen & Dolan, 2011a) and MSRVTT-QA (Xu et al., 2017). The results are presented in Table 6. Our model achieves the highest scores compared to all baselines. Due to the lack of video (frame sequence) generation benchmarks in our setting, we provide video generation examples in §3.5. These results demonstrate the superior performance of our models in both video understanding and video generation.
Table 7: Language-only evaluation. “I” denotes the instruction-tuned version.
Models MMLU
LLAMA-1-7B-Base 33.0
LLAMA-2-7B-Chat 47.9
SEED-LLAMA-8B-I 36.1
AnyGPT-Base 26.4
AnyGPT-Chat 27.4
MIO-Instruct 45.7
Table 8: Results for trimodal comprehension (text, image, and speech).
Models OmniBench
Gemini-1.5-Pro 42.67
Reka-Core-20240501 31.52
AnyGPT (8B) 17.77
video-SALMONN (13B) 34.11
Unified-IO 2 (6.8B) 34.24
MIO-Instruct (7B) 36.96
3.4 Language-only Tasks
We evaluate our models on MMLU (Hendrycks et al., 2021). The baselines are two LLaMA variants (Touvron et al., 2023a; b), the instruction-tuned SEED-LLaMA (Ge et al., 2023b), and AnyGPT (Zhan et al., 2024). For the MMLU benchmark, we conduct zero-shot evaluation experiments using the official evaluation code. The experimental results are shown in Table 8. We can observe that our models have superior language-only performance compared with all any-to-any MM-LLM baselines and even surpass LLaMA-1-7B-Base, an advanced pure language model.
@@ 3.5 Demonstrations
We illustrate the basic and advanced abilities of MIO in Figure 5 and 4. The basic abilities of MIO involve image understanding and generation, video understanding and generation, ASR, and TTS. The advanced abilities of MIO are based on its any-to-any and multimodal interleaved sequence generation features. These abilities involve visual storytelling (i.e., interleaved video-text generation), chain of visual thought, speech-in-speech-out, instructional image editing, visual guideline generation, etc. We refer the readers to Appendix E.5 for more demonstrations including multimodal chain of thought and multimodal in-context learning.
3.6 Ablation Studies
Generality for Trimodal Understanding.
We evaluate our model using the OmniBench (Li et al., 2024d), which incorporates text, image, and speech modalities as inputs, requiring the model to choose one of four options as the correct answer to determine accuracy. Although MIO acquires its multimodal understanding capabilities through dual-modal training, the evaluation results in Table 8 indicate that MIO also exhibits superior trimodal comprehension abilities.
Effect of Different Image Tokenizers.
The image tokenizer has a significant impact on image modality alignment. In Figure 2, we compare the image generation performance under a controlled setting after training for solely 3K steps in stage 1, using various image tokenizers. The image tokenizers used for comparison include a VQGAN (Esser et al., 2020) with a vocabulary size of 1024 and a compression rate of 16 (VQGAN-1024), as well as the VQGAN-Gumbel with a vocabulary size of 8192 (VQGAN-8192) [^5] https://github.com/CompVis/taming-transformers. Our results indicate that the SEED-Tokenizer, which captures more semantic and higher-level image information, exhibits faster convergence. In contrast, both VQGAN tokenizers show slower convergence due to their lower-level image information.
Figure 2: Comparing different image tokenizers for image generation within a controlled setting (limited to 3K training steps).
@@ 4.1 Multimodal LLMs
With the rapid success of Large Language Models (LLMs), current multimodal LLMs (MM-LLMs) are typically built on a pre-trained LLM backbone and are endowed with the ability to understand multiple modalities (Li et al., 2019; Lu et al., 2019; Kim et al., 2021; Zeng et al., 2022; Zhou et al., 2022; Wang et al., 2023b; 2024e). Generally, these MM-LLMs align the representations of images obtained from visual encoders with the text embedding space, thereby leveraging the powerful capabilities of the foundational models. For example, BLIP-2 (Li et al., 2023b) uses CLIP-ViT (Radford et al., 2021) to extract high-level features from images and then employs a Q-Former to compress the number of image tokens and further align image tokens with the LLM embeddings. In contrast, LLaVA (Liu et al., 2023b; Li et al., 2024a) utilizes a simple linear projection or MLP as the connector between the image encoder and the LLM backbone. These models demonstrate strong multimodal understanding abilities, achieving significant progress in tasks such as visual question answering, visual commonsense reasoning, chart understanding, etc.
Additionally, beyond images, other MM-LLMs have also focused on modalities such as speech and video. For instance, LLaSM (Shu et al., 2023) and InternVideo (Wang et al., 2022; 2024c) are MM-LLMs designed for speech and video understanding, respectively. These models adopt a similar architectural design to BLIP-2 or LLaVA but redesign modality-specific encoders.
Recently, increasing attention has been paid to unifying multiple modalities within a single MM-LLM. For example, ImageBind (Girdhar et al., 2023) develops encoders suited for multiple modalities such as images, videos, audio, heat maps, among others, while OmniBind (Wang et al., 2024d) trains an omni-representation model by aligning encoders across four modalities: audio, language, images, and 3D objects. OmniBench (Li et al., 2024d) is proposed to evaluate the models’ abilities for visual, acoustic, and textual understanding.
@@ 4.2 Any-to-Any MM-LLMs
To enable multimodal generation in MM-LLMs, a straightforward approach is to allow these models to call external multimodal generation tools, such as Stable Diffusion (Rombach et al., 2022) or text-to-speech (TTS) tools (Shen et al., 2023; Li et al., 2024c; OpenAI et al., 2023). However, as highlighted in the Gemini technical report (Team et al., 2023), relying on an intermediate natural language interface can limit the model’s ability to express images. If a model cannot natively output images, it will not be able to generate images with prompts of interleaved sequences of image and text. This claim is in line with our distinction between descriptive image generation and context-aware image generation, as discussed in §2.2.
As a result, recent works focus on the unification of multimodal understanding and generation in a single model (i.e., any-to-any MM-LLMs), enabling the generation of multimodal tokens without natural language as an interface. These models typically follow different approaches, depending on how images are represented in both input and output sides. For example, the Discrete-In-Discrete-Out (DIDO) approach has been explored in works such as SEED-LLaMA (Ge et al., 2023b), AnyGPT (Zhan et al., 2024), and Chameleon (Team, 2024). Continuous-In-Discrete-Out (CIDO) methods have been implemented in models like DaVinCi (Diao et al., 2023), Gemini (Team et al., 2023), and Unified-IO 2 (Lu et al., 2023). The Continuous-In-Continuous-Out (CICO) approach is used in models such as Emu (Sun et al., 2023c; a), and DreamLLM (Dong et al., 2023). Another approach, the integration of autoregression and diffusion (AR + Diff), can be seen in models like Transfusion (Zhou et al., 2024), Show-o (Xie et al., 2024), and Li et al. (2024b)’s.
However, these models face specific limitations. DreamLLM (CICI,  Dong et al. (2023)) and CIDO models suffer from inconsistencies between input and output forms for multimodal data, making it difficult for them to natively support the generation of interleaved multimodal sequences where an image functions in a coupled way as both input and output. Emu2 (CICO,  Sun et al. (2023a)) struggles with the challenges of the mean square error (MSE) loss used for training continuous output representations, as well as with the uni-modal assumption of the Gaussian distribution in the MSE loss. Transfusion (AR + Diff,  Zhou et al. (2024)) applies noise to images from the input side to support multimodal generation with diffusion modeling, and relies on VAE (Kingma & Welling, 2013) features rather than CLIP (Radford et al., 2021) features for denoising, which largely trade off the multimodal understanding abilities.
To mitigate these issues, we adopt the DIDO approach. A comprehensive comparison of our models with other any-to-any MM-LLMs is presented in Table 1.
@@ 5 Conclusion
Appendix A Pre-training Data
@@
    Image-text paired data: SBU (Ordonez et al., 2011), CC3M (Sharma et al., 2018), LAION-COCO (LAION, 2022), and JourneyDB (Pan et al., 2023), where JourneyDB only serves for image generation.
@@
    Language-only data: RefinedWeb (Penedo et al., 2023).
@@
    Image-text interleaved data: OBELICS (Laurençon et al., 2023), MMC4-core-ff (Zhu et al., 2023).
@@
    Video-text paired data: WebVid-10M (Bain et al., 2021).
@@
    Video-text interleaved data: HowTo-100M (Miech et al., 2019), Youtube-Temporal-180M (Zellers et al., 2021).
@@
    Speech-text paired data: Libriheavy (Kang et al., 2023).
@@
We have different data processing procedures for different data types illustrated in §A following Emu (Sun et al., 2023c) and Qwen-VL (Bai et al., 2023):
@@
    Language-only data: we use the same data processing pipeline as used in Yi (AI et al., 2024).
@@
    Image-text interleaved data: we filter the data using a CLIP score threshold of 0.25, and follow the same procedure as illustrated in Emu (Sun et al., 2023c).
@@
    Video-text interleaved data: We first use PySceneDetect to extract key frames from the video based on scene changes, following the practice of Stable Video Diffusion (Blattmann et al., 2023). Then, for each video clip between two key frames, we extract a central frame for textual caption generation with BLIP-2 (Li et al., 2023b). Additionally, the video clips between key frames are processed using ASR (automatic speech recognition) tools to extract subtitles. The ASR text and captions are then integrated and refined using Yi-34B-Chat (AI et al., 2024), resulting in a single text segment. These text segments, along with the key frames and central frames, form the video-text interleaved data.
@@
Appendix B Pre-training Details